OpenStudy (anonymous):

(a^2b)^8(a^3b^4) Simplify

OpenStudy (anonymous):

A^19B^12 I think...

OpenStudy (anonymous):

\[(a^{3}b^{1})^{8} \times (a^{3}b^{4}) = (a^{3\times8}\times b^{1\times 8}) \times a^{3}b^4\]\[= ...?\]